A railroad car with a mass of 12,000 kg collides with another with a mass of 20,000 kg that is at rest initially and they couple after the collision. If the car that comes to collide initially has a speed of 12 m/s, what is their speed together after the collision in m/s.


1
Expert's answer
2021-11-09T10:58:32-0500

We can find the common speed of the cars after the collision from the law of conservation of momentum:


"m_1v_1=(m_1+m_2)v_c,""v_c=\dfrac{m_1v_1}{m_1+m_2}=\dfrac{12000\ kg\times12\ \dfrac{m}{s}}{12000\ kg+20000\ kg}=4.5\ \dfrac{m}{s}."